Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A committee of 3 has to be formed randomly...

Expert replies
by BTGmoderatorLU » Sun Jan 07, 2018 7:55 am
A committee of 3 has to be formed randomly from a group of 6 people. If Tom and Mary are in this group of 6, what is the probability that Tom will be selected into the committee but Mary will not?

A. 1/10
B. 1/5
C. 3/10
D. 2/5
E. 1/2

The OA is C.

I'm really confused with this PS question. Experts, any suggestion about how can I solve it? Thanks in advance.
Join the discussion
Source: — Problem Solving |

by [email protected] » Sun Jan 07, 2018 4:41 pm
Hi LUANDATO,

We're told that a committee of 3 has to be formed randomly from a group of 6 people (including Tom and Mary). We're asked for the probability that Tom will be selected into the committee but Mary will not. This question can be solved by using the Combination Formula twice (in slightly different ways).

To start, we should calculate the total possible groups of 3... 6!/3!3! = (6)(5)(4)/(3)(2)(1) = 20 possible groups of 3

At this point, since there are so few possibilities, you could just list them all out. However, if you want to use Combination Formula to find the total number of groups that includes Tom but NOT Mary, here's how you would do it. We assign one of the 3 'spots' to Tom, meaning that there are 2 remaining spots (and neither of the two can go to Mary). This means that there are 4 possible people who can take those two spots... 4!/2!2! = (4)(3)/(2)(1) = 6 possible groups with Tom but NOT Mary

Probability of a group of 3 having Tom but NOT Mary = 6/20 = 3/10

Final Answer: C

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by GMATGuruNY » Mon Jan 08, 2018 3:39 am
LUANDATO wrote:A committee of 3 has to be formed randomly from a group of 6 people. If Tom and Mary are in this group of 6, what is the probability that Tom will be selected into the committee but Mary will not?

A. 1/10
B. 1/5
C. 3/10
D. 2/5
E. 1/2
From the 6 people, 3 will be selected.
Thus, P(Tom is selected) = 3/6 = 1/2.
From the 5 remaining people, 3 will NOT be selected.
Thus, P(Mary is not selected) = 3/5.
To combine these probabilities, we multiply:
1/2 * 3/5 = 3/10.

The correct answer is C.

Alternate approach:

ONE WAY:
One way to yield a favorable outcome is to choose Aaron first.
P(1st person selected is Aaron) = 1/6.
P(2nd person selected is not Karen) = 4/5. (Of the 5 remaining people, anyone but Karen.)
P(3rd person selected is not Karen) = 3/4. (Of the 4 remaining people, anyone but Karen.)
To combine these probabilities, we multiply:
1/6 * 4/5 * 3/4 = 1/10.

OTHER WAYS:
Since a favorable outcome will be achieved if Aaron is chosen 1st, 2nd, or 3rd -- for a total of 3 ways -- we multiply by 3:
1/10 * 3 = 3/10.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Scott@TargetTestPrep » Sun Aug 04, 2019 10:30 am
BTGmoderatorLU wrote:A committee of 3 has to be formed randomly from a group of 6 people. If Tom and Mary are in this group of 6, what is the probability that Tom will be selected into the committee but Mary will not?

A. 1/10
B. 1/5
C. 3/10
D. 2/5
E. 1/2

The OA is C.

I'm really confused with this PS question. Experts, any suggestion about how can I solve it? Thanks in advance.
The total number of committees of 3 people from a group of 6 people that can be formed is 6C3 = (6 x 5 x 4)/(3 x 2) = 20. The number of committees can be formed in which Tom will be selected and Mary will not consists of Tom’s being chosen for a spot, and 2 of 4 individuals (but not Mary) chosen for the remaining 2 spots:

1C1 x 1C0 x 4C2 = 1 x 1 x (4 x 3)/2 = 6

Therefore, the probability that Tom will be selected for the committee but Mary will not is 6/20 = 3/10.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion